### v1

# A 3.3904-Competitive Online Algorithm for List Update with Uniform Costs

Abstract
We consider the List Update problem where the cost of each swap is assumed to be 1. This is in contrast to the “standard” model, in which an algorithm is allowed to swap the requested item with previous items for free. We construct an online algorithm Full-Or-Partial-Move (Fpm), whose competitive ratio is $R\approx 3.3904$, improving over the previous best known bound of $4$.
keywords:
List update, online algorithms, competitive analysis
1 Introduction
The List Update problem.
In the online List Update problem, the objective is to maintain a set of items stored in a linear list in response to a sequence of access requests. The cost of accessing a requested item is equal to its distance from the front of the list. After each request, an algorithm is allowed to rearrange the list by executing any number of swaps of adjacent items. In the model introduced by Sleator and Tarjan in their seminal 1985 paper on competitive analysis [25], an algorithm can repeatedly swap the requested item with its preceding item at no cost. These swaps are called free. All other swaps are called paid and have cost $1$ each. As in other problems involving self-organizing data structures (see [7]), the objective of an algorithm is to judiciously adjust the list at each step to assure that, independently of future requests, its total cost is never too big in comparison to the optimum cost. If an algorithm’s cost is always within a factor $c$ of the optimum, it is said to be $c$-competitive, and this constant $c$ is referred to as its competitive ratio.
Sleator and Tarjan proved that Algorithm Move-to-Front (Mtf), that after each request moves the requested item to the front of the list, is $2$-competitive. This ratio is known to be optimal if the number of items is unbounded. Their work culminated prior extensive studies of List Update, that included experimental results, probabilistic approaches, and earlier attempts at amortized analysis (see [14] and the references therein).
As shown in subsequent work, Mtf is not unique — there are other strategies that achieve ratio $2$. One example is algorithm TimeStamp from [2]. Anderson et al. [9] showed that there is a large class of $2$-competitive algorithms based on work functions. In fact, it is known that there are infinitely many algorithms that achieve ratio $2$ [12, 21].
The uniform cost model.
Following [24], we will refer to the cost model of [25] as standard, and we will denote it here by $\textsf{LUP}_{0}$. This model has been questioned in the literature for not accurately reflecting true costs in some implementations [24, 23, 22, 18], with the concept of free swaps being one of the main concerns.
A natural approach to address this concern, considered in some later studies of the List Update problem, is to simply charge cost $1$ for any swap. (See, for example, [24, 4, 11].) We will call it here the uniform cost model and denote it $\textsf{LUP}_{1}$. Changing the cost of free swaps from 00 to $1$ at most doubles any algorithm’s cost, so Mtf is no worse than $4$-competitive for $\textsf{LUP}_{1}$, and it is also easy to show that this bound is tight. A general lower bound of $3$ on the competitive ratio of deterministic algorithms for $\textsf{LUP}_{1}$ was given by Reingold et al. [24]. Natural adaptations of $2$-competitive algorithms for $\textsf{LUP}_{0}$ can be shown to have ratio larger than $3$ (see the discussion in section 5), and no algorithm is known to beat ratio $4$. In summary, in contrast to the standard model, closing or reducing this $3$-to-$4$ gap between the lower and upper bound for $\textsf{LUP}_{1}$ remains elusive [^1] The authors of [21] claimed an upper bound of $3$ for $\textsf{LUP}_{1}$, but discovered later that their proof was not correct (personal communication with S. Kamali)..
Our results.
Addressing this open problem, we develop an online algorithm for $\textsf{LUP}_{1}$ with competitive ratio ${\frac{1}{8}}\!\cdot\!(23+\sqrt{17})\approx 3.3904$, significantly improving the previous upper bound of $4$. Our algorithm, called Fpm, is based on a combination of several ideas. One involves the movement of the requested item. It can be shown (see section 5) that any algorithm that chooses either not to move at all or to move to front cannot be $3$-competitive. For this reason, Fpm will sometimes make a partial move to a suitably chosen target location in the list. Another idea involves keeping track, for each pair of items, of the work function for the two-item subsequence consisting of these items. These work functions are used in two ways. On one hand, they indicate, roughly, which relative order between the items in each pair is more likely in an optimal solution. The algorithm uses this information to determine the target location of each item. On the other hand, the simple sum of all these pair-based work functions is a lower bound on the optimum cost, which comes in handy in the competitive analysis of Fpm.
Fpm remains $3.3904$-competitive even for the partial cost function, where the cost to access location $\ell$ is $\ell-1$, instead of the full cost of $\ell$ used in the original definition of List Update in [25]. Both functions have been alternatively used in the literature, depending on context and convenience. For any online algorithm, its competitive ratio for partial costs is at least as large as for full costs, although the difference typically vanishes if the list size is unbounded. We present our analysis in terms of partial costs. Further, the $3.3904$-competitiveness proof for Fpm easily extends to the dynamic scenario of List Update that also allows operations of insertions and deletions as in the original definition from [25].
The most intriguing question left open in our work is whether competitive ratio $3$ can be achieved. We discuss it in section 5. As explained there, we verified computationally that $3$-competitive algorithms exist for lists with up to $6$ items. On the other hand, several natural strategies (including our Fpm) fail to produce a general $3$-competitive algorithm.
Other related work.
Randomization is a natural tool in the online setting, and for many online problems it can significantly improve the competitive ratio. For $\textsf{LUP}_{0}$, the study of randomized algorithms was initiated by Irani [19] who showed an upper bound of $\approx 1.94$, the first bound below $2$. This was subsequently improved in [24, 2, 6], with the last paper, by Albers et al., giving the best currently known upper bound of $1.6$. This is already quite close to the lower bound of $1.5$, given by Teia [26] in 1993. In fact, in the partial-cost model there is a matching lower bound [8] for so-called projective algorithms, whose computation is uniquely determined by their behavior on two-item instances. For the uniform cost model ($\textsf{LUP}_{1}$), Reingold et al. [24] show an upper bound of $\approx 2.64$. Recent work by Albers and Janke [4] implies that this ratio can be improved to $\approx 2.4$.
Other cost models have been investigated. In the earlier mentioned papers [24, 4] on randomized algorithms, the authors generalize the uniform cost function by distinguishing between the cost of $1$ for following a link during search and the cost of $d\geq 1$ for a swap. A completely different approach is discussed in [22, 18], where an algorithm is allowed to completely rearrange the list at the cost that is only linear with respect to the position of the accessed item. Yet another approach is discussed in [23]. In these last two models no algorithm can be constant-competitive.
Speaking to the centrality of the List Update problem within the area of competitive analysis, a variety of its variants has been investigated in the literature over the last forty years, including models with lookahead [1], locality of reference [10, 5], parametrized approach [17] and algorithms with advice [15]. A particularly interesting model was proposed recently by Azar et al. [11], where an online algorithm is allowed to postpone serving some requests, but is either required to serve it by a specified deadline or has to pay some delay penalty.
There is also a number of survey papers on the List Update problem, where the reader can find more information about this problem and its variants [3, 20, 21].
2 Preliminaries
Model.
An algorithm has to maintain a list of items, while a sequence $\sigma$ of access requests is presented to an algorithm in an online manner. In each step $t\geq 1$ an algorithm is presented an access request $\sigma^{t}$ to an item in the list. If this item is in a location $\ell$, the algorithm incurs cost $\ell-1$ to access it. (The locations in the list are indexed $1,2,...$.) Afterwards, the algorithm may change the list configuration by performing an arbitrary number of swaps of neighboring items, each of cost $1$.
For any algorithm $A$, we denote its cost for processing a sequence $\sigma$ by $A(\sigma)$. Furthermore, we denote the optimal algorithm by Opt.
Notation.
Let $\mathcal{P}$ be the set of all unordered pairs of items. For a pair $\{x,y\}\in\mathcal{P}$, we use the notation $x\prec y$ ($x\succ y$) to denote that $x$ is before (after) $y$ in the list of an online algorithm. (The relative order of $x,y$ may change over time, but it will be always clear from context what step of the computation we are referring to.) We use $x\preceq y$ ($x\succeq y$) to denote that $x\prec y$ ($x\succ y$) or $x=y$.
For an input $\sigma$ and a pair $\{x,y\}\in\mathcal{P}$, $\sigma_{xy}$ is the subsequence of $\sigma$ restricted to requests to items $x$ and $y$ only. Whenever we say that an algorithm serves input $\sigma_{xy}$, we mean that an algorithm has to maintain a list of two items, $x$ and $y$.
2.1 Work Functions
Work functions on item pairs.
For each prefix $\sigma$ of the input sequence, an online algorithm may compute a so-called work function $W^{xy}$, where $W^{xy}(xy)$ (or $W^{xy}(yx)$) is the optimal cost of the solution that serves $\sigma_{xy}$ and ends with the list in configuration $xy$ (or $yx$). (Function $W^{xy}$ also has prefix $\sigma$ as an argument. Its value will be always uniquely determined from context.) The values of $W$ for each step can be computed iteratively using straightforward dynamic programming. Note that the values of $W$ are non-negative integers and $|W^{xy}(xy)-W^{xy}(yx)|\leq 1$.
Modes.
For a pair $\{x,y\}\in\mathcal{P}$, we define its mode depending on the value of the work function $W^{xy}$ in the current step and the mutual relation of $x$ and $y$ in the list of an online algorithm. In the following definition we assume that $y\prec x$.
  • Pair $\{x,y\}$ is in mode $\alpha$ if $W^{xy}(yx)+1=W^{xy}(xy)$.
  • Pair $\{x,y\}$ is in mode $\beta$ if $W^{xy}(yx)=W^{xy}(xy)$.
  • Pair $\{x,y\}$ is in mode $\gamma$ if $W^{xy}(yx)-1=W^{xy}(xy)$.
For an illustration of work function evolution and associated modes, see Figure 1.
Figure 1: The evolution of the work function $W^{xy}$. Initially, $W^{xy}$ has its minimum in the state $yx$ and $y\prec x$ (in the list of an online algorithm). Thus, the mode of the pair $\{x,y\}$ is $\alpha$. Next, because of the two requests to $x$, the value of $W^{xy}(yx)$ is incremented, while the value at $xy$ remains intact. The mode is thus changed from $\alpha$ to $\beta$ and then to $\gamma$. Finally, when an algorithm moves item $x$ before $y$, the mode of the pair $\{x,y\}$ changes to $\alpha$. The value of $w^{xy}$ (the average of $W^{xy}(xy)$ and $W^{xy}(yx)$) increases by ${\frac{1}{2}}$ whenever the mode changes due to a request.
If a pair $\{x,y\}$ is in mode $\alpha$, then the minimum of the work function $W^{xy}$ is at configuration $yx$, i.e., the one that has $y$ before $x$. That is, an online algorithm keeps these items in a way that “agrees” with the work function. Note that $\alpha$ is the initial mode of all pairs. Conversely, if a pair $\{x,y\}$ is in state $\gamma$, then the minimum of the work function $W^{xy}$ is at configuration $xy$. In this case, an online algorithm keeps these items in a way that “disagrees” with the work function.
2.2 Lower Bound on OPT
Now we show how to use the changes in the work functions of item pairs to provide a useful lower bound on the cost of an optimal algorithm. The following lemma is a standard and straightforward result of the list partitioning technique [9]. [^2] There are known input sequences on which the relation of 2.1 is not tight [9].
Lemma 2.1.
For every input sequence $\sigma$, it holds that $\sum_{\{x,y\}\in\mathcal{P}}\textsl{{Opt}}(\sigma_{xy})\leq\textsl{{Opt}}(\sigma)$.
Averaging work functions.
We define the function $w^{xy}$ as the average value of the work function $W^{xy}$, i.e.,
$w^{xy}\triangleq\textstyle{\frac{1}{2}}\cdot(\,W^{xy}(xy)+W^{xy}(yx)\,).$
We use $w_{t}^{xy}$ to denote the value of $w^{xy}$ after serving the first $t$ requests of $\sigma$, and define $\Delta_{t}w^{xy}\triangleq w_{t}^{xy}-w_{t-1}^{xy}$. The growth of $w^{xy}$ can be related to $\textsl{{Opt}}(\sigma)$ in the following way.
Lemma 2.2.
For a sequence $\sigma$ consisting of $T$ requests, it holds that $\sum_{t=1}^{T}\sum_{\{x,y\}\in\mathcal{P}}\Delta_{t}w^{xy}\leq\textsl{{Opt}}(\sigma).$
Proof 2.3.
We first fix a pair $\{x,y\}\in\mathcal{P}$. We have $w_{0}^{xy}=\textstyle{\frac{1}{2}}$ and $w_{T}^{xy}\leq\textsl{{Opt}}(\sigma_{xy})+\textstyle{\frac{1}{2}}$. Hence, $\sum_{t=1}^{T}\Delta_{t}w^{xy}=w_{T}^{xy}-w_{0}^{xy}\leq\textsl{{Opt}}(\sigma_{xy}).$ The proof follows by summing over all pairs $\{x,y\}\in\mathcal{P}$ and invoking 2.1.
Pair-based OPT.
2.2 gives us a convenient tool to lower bound $\textsl{{Opt}}(\sigma)$. We define the cost of pair-based Opt in step $t$ as $\sum_{\{x,y\}\in\mathcal{P}}\Delta_{t}w^{xy}$. For a given request sequence $\sigma$, the sum of these costs over all steps when processing $\sigma$ is a lower-bounds on the actual value of $\textsl{{Opt}}(\sigma)$.
On the other hand, we can express $\Delta_{t}w^{xy}$ (and thus also the pair-based Opt) in terms of the changes of the modes of item pairs. See Figure 1 for an illustration.
{observation}
If a pair $\{x,y\}$ changes its mode due to the request in step $t$ then $\Delta_{t}w^{xy}=\textstyle{\frac{1}{2}}$, otherwise $\Delta_{t}w^{xy}=0$.
3 Algorithm Full-Or-Partial-Move
For each item $x$, Fpm keeps track of an item denoted $\theta_{x}$ and called the target of $x$. At the beginning, Fpm sets $\theta_{x}=x$ for all $x$. Furthermore, at each time, Fpm ensures that $\theta_{x}\preceq x$.
The rest of this section describes the overall strategy of algorithm Fpm. Our description is top-down, and proceeds in three steps:
  • First we describe, in broad terms, what actions are involved in serving a request, including the choice of a move and the principle behind updating target items.
  • Next, we define the concept of states associated with item pairs and their potentials.
  • Finally, we explain how algorithm Fpm uses these potential values to decide how to adjust the list after serving the request.
This description will fully specify how Fpm works, providing that the potential function on the states is given. Thus, for any choice of the potential function the competitive ratio of Fpm is well-defined. What remains is to choose these potential values to optimize the competitive ratio. This is accomplished by the analysis in section 4 that follows.
Serving a request.
Whenever an item $z^{*}$ is requested, Fpm performs the following three operations, in this order:
1. Target cleanup. If $z^{*}$ was a target of another item $y$ (i.e., $\theta_{y}=z^{*}$ for $y\neq z^{*}$), then $\theta_{y}$ is updated to the successor of $z^{*}$. This happens for all items $y$ with this property.
2. Movement of $z^{*}$. Fpm executes one of the two actions: a partial move or a full move. We will explain how to choose between them later.
  • In the partial move, item $z^{*}$ is inserted right before $\theta_{z^{*}}$. (If $\theta_{z^{*}}=z^{*}$, this means that $z^{*}$ does not change its position.)
  • In the full move, item $z^{*}$ is moved to the front of the list.
3. Target reset. $\theta_{z^{*}}$ is set to the front item of the list.
It is illustrative to note a few properties and corner cases of the algorithm.
  • Target cleanup is executed only for items following $z^{*}$, and thus the successor of $z^{*}$ exists then (i.e., Fpm is well defined).
  • If $\theta_{z^{*}}=z^{*}$ and a partial move is executed, then $z^{*}$ is not moved, but the items that targeted $z^{*}$ now target the successor of $z^{*}$.
  • For an item $x$, the items that precede $\theta_{x}$ in the list were requested (each at least once) after the last time $x$ had been requested.
Modes, flavors and states.
Fix a pair $\{x,y\}$ such that $y\prec x$, and thus also $\theta_{y}\preceq y\prec x$. This pair is assigned one of four possible flavors, depending on the position of $\theta_{x}$ (cf. Figure 2):
  • flavor $d$ (disjoint): if $\theta_{y}\preceq y\prec\theta_{x}\preceq x$,
  • flavor $o$ (overlapping): if $\theta_{y}\prec\theta_{x}\preceq y\prec x$,
  • flavor $e$ (equal): if $\theta_{y}=\theta_{x}\preceq y\prec x$,
  • flavor $n$ (nested): if $\theta_{x}\prec\theta_{y}\preceq y\prec x$
Figure 2: The flavor of a pair $\{x,y\}$ (where $y\prec x$) depends on the position of target $\theta_{x}$ with respect to items $\theta_{y}\preceq y\prec x$. In the figure for flavor $e$, we write $\theta_{xy}$ for the item $\theta_{x}=\theta_{y}$.
For a pair $\{x,y\}$ that is in a mode $\xi\in\{\alpha,\beta,\gamma\}$ and has a flavor $\omega\in\{d,o,e,n\}$, we say that the state of $\{x,y\}$ is $\xi^{\omega}$. Recall that all pairs are initially in mode $\alpha$, and note that their initial flavor is $d$. That is, the initial state of all pairs is $\alpha^{d}$.
We will sometimes combine the flavors $o$ with $e$, stating that the pair is in state $\xi^{oe}$ if its mode is $\xi$ and its flavor is $o$ or $e$. Similarly, we will also combine flavors $n$ and $e$.
Pair potential.
To each pair $\{x,y\}$ of items, we assign a non-negative pair potential $\Phi_{xy}$. We abuse the notation and use $\xi^{\omega}$ not only to denote the pair state, but also the corresponding values of the pair potential. That is, we assign the potential $\Phi_{xy}=\xi^{\omega}$ if pair $\{x,y\}$ is in state $\xi^{\omega}$. We pick the actual values of these potentials only later in subsection 4.5.
We emphasize that the states of each item pair depend on work functions for this pair that are easily computable in online manner. Thus, Fpm may compute the current values of $\Phi_{xy}$, and also compute how their values would change for particular choice of a move.
Choosing the cheaper move.
For a step $t$, let $\Delta_{t}\textsl{{Fpm}}$ be the cost of Fpm in this step, and for a pair $\{x,y\}\in\mathcal{P}$, let $\Delta_{t}\Phi_{xy}$ be the change of the potential of pair $\{x,y\}$ in step $t$. Let $z^{*}$ denote the requested item. Fpm chooses the move (full or partial) with the smaller value of
$\Delta_{t}\textsl{{Fpm}}+\sum_{y\prec z^{*}}\Delta_{t}\Phi_{z^{*}y},$
breaking ties arbitrarily. Importantly, note that the value that Fpm minimizes involves only pairs including the requested item $z^{*}$ and items currently preceding it.
4 Analysis of Full-Or-Partial-Move
In this section, we show that for a suitable choice of pair potentials, Fpm is $3.3904$-competitive.
To this end, we first make a few observations concerning how the modes, flavors (and thus states) of item pairs change due to the particular actions of Fpm. These are summarized in Table 1, Table 2, and Table 3, and proved in subsection 4.2 and subsection 4.3.
Next, in subsection 4.4 and subsection 4.5, we show how these changes influence the amortized cost of Fpm pertaining to particular pairs. We show that for a suitable choice of pair potentials, we can directly compare the amortized cost of Fpm with the cost of Opt.
4.1 Structural Properties
Note that requests to items other than $x$ and $y$ do not affect the mutual relation between $x$ and $y$. Moreover, to some extent, they preserve the relation between targets $\theta_{x}$ and $\theta_{y}$, as stated in the following lemma.
Lemma 4.1.
Fix a pair $\{x,y\}\in\mathcal{P}$ and assume that $\theta_{y}\preceq\theta_{x}$ (resp. $\theta_{y}=\theta_{x}$). If Fpm serves a request to an item $z^{*}$ different than $x$ and $y$, then these relations are preserved. The relation $\theta_{y}\prec\theta_{x}$ changes into $\theta_{y}=\theta_{x}$ only when $\theta_{y}$ and $\theta_{x}$ are adjacent and $z^{*}=\theta_{y}$.
Proof 4.2.
By the definition of Fpm, the targets of non-requested items are updated (during the target cleanup) only if they are equal to $z^{*}$. In such a case, they are updated to the successor of $z^{*}$. We consider several cases.
  • If $z^{*}\notin\{\theta_{x},\theta_{y}\}$, the targets $\theta_{x}$ and $\theta_{y}$ are not updated.
  • If $z^{*}=\theta_{y}=\theta_{x}$, both targets are updated to the successor of $z^{*}$, and thus they remain equal.
  • If $z^{*}=\theta_{y}\prec\theta_{x}$, target $\theta_{y}$ is updated. If $\theta_{y}$ and $\theta_{x}$ were adjacent ($\theta_{y}$ was an immediate predecessor of $\theta_{x}$), they become equal. In either case, $\theta_{y}\preceq\theta_{x}$.
  • If $\theta_{y}\prec\theta_{x}=z^{*}$, target $\theta_{x}$ is updated, in which case the relation $\theta_{y}\prec\theta_{x}$ is preserved.
4.2 Mode Transitions
Now, we focus on the changes of modes. It is convenient to look first at how they are affected by the request itself (which induces an update of the work function), and subsequently due to the actions of Fpm (when some items are swapped). The changes are summarized in the following observation.
{observation}
Let $z^{*}$ be the requested item.
  • Fix an item $y\succ z^{*}$. The mode transitions of the pair $\{z^{*},y\}$ due to request are $\alpha\to\alpha$, $\beta\to\alpha$, and $\gamma\to\beta$. Subsequent movement of $z^{*}$ does not further change the mode.
  • Fix an item $y\prec z^{*}$. Due to the request, pair $\{z^{*},y\}$ changes first its mode due to the request in the following way: $\alpha\to\beta$, $\beta\to\gamma$, and $\gamma\to\gamma$. Afterwards, if Fpm moves $z^{*}$ before $y$, the subsequent mode transitions for pair $\{z^{*},y\}$ are $\beta\to\beta$ and $\gamma\to\alpha$.
4.3 State Transitions
Throughout this section, we fix a step and let $z^{*}$ be the requested item.
We analyze the potential changes for both types of movements. We split our considerations into three cases corresponding to three types of item pairs. The first two types involve $z^{*}$ as one pair item, where the second item either initially precedes $z^{*}$ (cf. 4.5) or follows $z^{*}$ (cf. 4.7). The third type involves pairs that do not contain $z^{*}$ at all (cf. 4.9).
While we defined 12 possible states (3 modes $\times$ 4 flavors), we will show that $\alpha^{n}$, $\gamma^{d}$, and $\gamma^{o}$ never occur. This clearly holds at the very beginning, as all pairs are then in state $\alpha^{d}$.
For succinctness, we also combine some of the remaining states, reducing the number of states to the following six: $\alpha^{d}$, $\beta^{d}$, $\alpha^{oe}$, $\beta^{o}$, $\beta^{ne}$ and $\gamma^{ne}$. (For example, a pair is in state $\alpha^{oe}$ if it’s in state $\alpha^{o}$ or $\alpha^{e}$.) In the following we analyze the transitions between them. We start with a simple observation.
Lemma 4.3.
Fix an item $y\neq z^{*}$. If Fpm performs a full move, then the resulting flavor of the pair $\{z^{*},y\}$ is $d$.
Proof 4.4.
Item $z^{*}$ is moved to the front of the list, and its target $\theta_{z^{*}}$ is reset to this item, i.e., $z^{*}=\theta_{z^{*}}$. After the movement, we have $z^{*}\prec\theta_{y}$. This relation follows trivially if $z^{*}$ is indeed moved. However, it holds also if $z^{*}$ was already on the first position: even if $\theta_{y}=z^{*}$ before the move, $\theta_{y}$ would be updated to the successor of $z^{*}$ during the target cleanup. The resulting ordering is thus $\theta_{z^{*}}=z^{*}\prec\theta_{y}\preceq y$, i.e., the flavor of the pair becomes $d$.
Lemma 4.5.
Fix $y\prec z^{*}$. The state transitions for pair $\{z^{*},y\}$ are given in Table 1.
Proof 4.6.
First, suppose Fpm performs a full move. The pair $\{z^{*},y\}$ is swapped, which changes its mode according to subsection 4.2 ($\alpha\to\beta$, $\beta\to\alpha$, $\gamma\to\alpha$). By 4.3, the flavor of the pair becomes $d$. This proves the correctness of the transitions in row three of Table 1.
In the rest of the proof, we analyze the case when Fpm performs a partial move. We consider three sub-cases depending on the initial flavor of the pair $\{z^{*},y\}$. For the analysis of mode changes we will apply subsection 4.2.
  • Before the movement, the flavor of the pair was $d$, i.e., $\theta_{y}\preceq y\prec\theta_{z^{*}}\preceq z^{*}$.
    The movement of $z^{*}$ does not swap the pair, i.e., its mode transitions are $\alpha\to\beta$ and $\beta\to\gamma$. As $\theta_{z^{*}}$ is set to the list front, after the movement $\theta_{z^{*}}\preceq\theta_{y}\preceq y\prec z^{*}$, i.e., the flavor of the pair becomes either $e$ or $n$. This explains the state transitions $\alpha^{d}\to\beta^{ne}$ and $\beta^{d}\to\gamma^{ne}$.
  • Before the movement, the flavor of the pair was $o$, i.e., $\theta_{y}\prec\theta_{z^{*}}\preceq y\prec z^{*}$.
    Due to the movement, the pair is swapped, and its mode transitions are $\alpha\to\beta$ and $\beta\to\alpha$. As $\theta_{z^{*}}$ is set to the list front, we have $\theta_{z^{*}}\preceq\theta_{y}$. After the movement, $\theta_{y}\preceq z^{*}\prec y$, i.e., the flavor of the pair becomes either $o$ or $e$. This explains the state transitions $\alpha^{oe}\to(\beta^{o}$ or $\beta^{ne})$ and $\beta^{o}\to\alpha^{oe}$.
  • Before the movement, the flavor of the pair was $e$ or $n$, i.e., $\theta_{z^{*}}\preceq\theta_{y}\preceq y\prec z^{*}$.
    The movement swaps the pair, and its mode transitions are $\alpha\to\beta$, $\beta\to\alpha$ and $\gamma\to\alpha$. After the movement $z^{*}\prec\theta_{y}$, and target $\theta_{z^{*}}$ is set to the list front, which results in $\theta_{z^{*}}\preceq z^{*}\prec\theta_{y}\preceq y$, i.e., the pair flavor becomes $d$. This explains the state transitions $\alpha^{oe}\to\beta^{d}$, $\beta^{ne}\to\alpha^{d}$, and $\gamma^{ne}\to\alpha^{d}$
State before move $\alpha^{d}$ $\beta^{d}$ $\alpha^{oe}$ $\beta^{o}$ $\beta^{ne}$ $\gamma^{ne}$
State after partial move $\beta^{ne}$ $\gamma^{ne}$ $\beta^{d}$ or $\beta^{o}$ or $\beta^{ne}$ $\alpha^{oe}$ $\alpha^{d}$ $\alpha^{d}$
State after full move $\beta^{d}$ $\alpha^{d}$ $\beta^{d}$ $\alpha^{d}$ $\alpha^{d}$ $\alpha^{d}$
Table 1: State transitions for pairs $\{z^{*},y\}$ where $y\prec z^{*}$ right before the request to $z^{*}$.
Lemma 4.7.
Fix $y\succ z^{*}$. The state transitions for pair $\{z^{*},y\}$ are given in Table 2.
State before move $\alpha^{d}$ $\beta^{d}$ $\alpha^{oe}$ $\beta^{o}$ $\beta^{ne}$ $\gamma^{ne}$
State after move $\alpha^{d}$ $\alpha^{d}$ $\alpha^{d}$ $\alpha^{d}$ $\alpha^{d}$ or $\alpha^{oe}$ $\beta^{d}$ or $\beta^{o}$ or $\beta^{ne}$
Table 2: State transitions for pairs $\{z^{*},y\}$ where $z^{*}\prec y$ right before the request to $z^{*}$.
Proof 4.8.
The pair $\{z^{*},y\}$ is not swapped due to the request, and thus, by subsection 4.2, the mode transition are $\alpha\to\alpha$, $\beta\to\alpha$, $\gamma\to\beta$.
If Fpm performs a full move, the flavor of the pair becomes $d$ by 4.3. This explains the state transitions $\alpha^{d}\to\alpha^{d}$, $\beta^{d}\to\alpha^{d}$, $\alpha^{oe}\to\alpha^{d}$, $\beta^{o}\to\alpha^{d}$, $\beta^{ne}\to\alpha^{d}$, and $\gamma^{ne}\to\beta^{d}$.
In the following, we assume that Fpm performs a partial move, and we will identify cases where the resulting pair flavor is different than $d$. We consider two cases.
  • The initial flavor is $d$, $o$ or $e$. That is $\theta_{z^{*}}\preceq z^{*}\prec y$ and $\theta_{z^{*}}\preceq\theta_{y}\preceq y$. During the target cleanup, $\theta_{y}$ may be updated to its successor, but it does not affect these relations, and in particular we still have $\theta_{z^{*}}\preceq\theta_{y}$. Thus, when $z^{*}$ is moved, it gets placed before $\theta_{y}$. This results in the ordering $\theta_{z^{*}}\preceq z^{*}\prec\theta_{y}\preceq y$, i.e., the resulting flavor is $d$.
  • The initial flavor is $n$, i.e., $\theta_{y}\prec\theta_{z^{*}}\preceq z^{*}\prec y$. As $\theta_{y}\neq z^{*}$, the target $\theta_{y}$ is not updated during the target cleanup. As $z^{*}$ is moved right before original position of $\theta_{z^{*}}$, it is placed after $\theta_{y}$, and the resulting ordering is $\theta_{z^{*}}\preceq\theta_{y}\prec z^{*}\prec y$. That is, the flavor becomes $o$ or $e$, which explains the state transitions $\beta^{ne}\to\alpha^{oe}$ and $\gamma^{ne}\to(\beta^{o}$ or $\beta^{ne})$.
State before move $\alpha^{d}$ $\beta^{d}$ $\alpha^{oe}$ $\beta^{o}$ $\beta^{ne}$ $\gamma^{ne}$
State after move $\alpha^{d}$ $\beta^{d}$ $\alpha^{oe}$ $\beta^{o}$ or $\beta^{ne}$ $\beta^{ne}$ $\gamma^{ne}$
Table 3: State transitions for pairs $\{x,y\}$ where $x\neq z^{*}$ and $y\neq z^{*}$ right before the request to $z^{*}$.
Lemma 4.9.
Fix $y\prec x$, such that $x\neq z^{*}$ and $y\neq z^{*}$. State transitions for pair $\{x,y\}$ are given in Table 3.
Proof 4.10.
The mode of the pair $\{x,y\}$ is not affected by the request to $z^{*}$.
The flavor of the pair $\{x,y\}$ depends on mutual relations between $x$, $y$, $\theta_{x}$ and $\theta_{y}$. By 4.1, the only possible change is that $\theta_{x}$ and $\theta_{y}$ were different but may become equal: this happens when they were adjacent and the earlier of them is equal to $z^{*}$. We consider four cases depending on the initial flavor of the pair.
  • The initial flavor was $e$ ($\theta_{y}=\theta_{x}\preceq y\prec x$). As $\theta_{y}$ and $\theta_{x}$ are not adjacent, the flavor remains $e$.
  • The initial flavor was $d$ ($\theta_{y}\preceq y\prec\theta_{x}\preceq x$). Suppose $\theta_{y}=z^{*}$. As $y\neq z^{*}$, we have $\theta_{y}=z^{*}\prec y\prec\theta_{x}$. That is, $\theta_{x}$ and $\theta_{y}$ are not adjacent, and thus the flavor remains $d$.
  • The initial flavor was $o$ ($\theta_{y}\prec\theta_{x}\preceq y\prec x$). In this case it is possible that $\theta_{y}$ and $\theta_{x}$ are adjacent and $z^{*}=\theta_{y}$. The flavor may thus change to $e$ or remain $o$.
  • The initial flavor was $n$ ($\theta_{x}\prec\theta_{y}\preceq y\prec x$). Similarly to the previous case, it is possible that $\theta_{x}$ and $\theta_{y}$ are adjacent and $z^{*}=\theta_{x}$. The flavor may thus change to $e$ or remain $n$.
4.4 Amortized Analysis
We set $R=\textstyle{\frac{1}{8}}(23+\sqrt{17})\leq 3.3904$ as our desired competitive ratio.
In the following, we fix a step $t$ in which item $z^{*}$ is requested. We partition $\mathcal{P}$ into three sets corresponding to the three types of pairs:
  • $\mathcal{P}^{t}_{1}\triangleq\{\{y,z^{*}\}:y\prec z^{*}\}$ (pairs where $z^{*}$ is the second item, analyzed in Table 1),
  • $\mathcal{P}^{t}_{2}\triangleq\{\{y,z^{*}\}:z^{*}\prec y\}$ (pairs where $z^{*}$ is the first item, analyzed in Table 2),
  • $\mathcal{P}^{t}_{3}\triangleq\{\{x,y\}:x\neq z^{*}\wedge y\neq z^{*}\}$ (pairs where $z^{*}$ is not involved, analyzed in Table 3).
For succinctness, wherever it does not lead to ambiguity, we omit subscripts $t$, i.e., write $\Delta\Phi_{xy}$ and $\Delta w^{xy}$ instead of $\Delta_{t}\Phi_{xy}$ and $\Delta_{t}w^{xy}$. We also omit superscripts $t$ in $P^{t}_{1}$, $P^{t}_{2}$, and $P^{t}_{3}$.
Our goal is to show the following three bounds:
  • $\Delta\textsl{{Fpm}}+\sum_{\{x,y\}\in\mathcal{P}_{1}}\Delta\Phi_{xy}\leq R\cdot\sum_{\{x,y\}\in\mathcal{P}_{1}}\Delta w^{xy}$,
  • $\sum_{\{x,y\}\in\mathcal{P}_{2}}\Delta\Phi_{xy}\leq R\cdot\sum_{\{x,y\}\in\mathcal{P}_{2}}\Delta w^{xy}$,
  • $\sum_{\{x,y\}\in\mathcal{P}_{3}}\Delta\Phi_{xy}\leq R\cdot\sum_{\{x,y\}\in\mathcal{P}_{3}}\Delta w^{xy}$.
Note that the left hand sides of these inequalities correspond to the portions of amortized cost of Fpm corresponding to sets $\mathcal{P}_{1},\mathcal{P}_{2},\mathcal{P}_{3}$, while the right hand sides are equal to $R$ times the corresponding portion of the cost of pair-based Opt. Hence, if we can show the above inequalities for every step $t$, the competitive ratio of $R$ will follow by simply adding them up.
As we show in the sections that follow, these bounds reduce to some constraints involving state potentials $\alpha^{d}$, $\beta^{d}$, $\alpha^{oe}$, $\beta^{o}$, $\beta^{ne}$ and $\gamma^{ne}$. The bounds for $\mathcal{P}_{2}$ and $\mathcal{P}_{3}$, while they involve summations over pairs, can be justified by considering individual pairs and the needed constraints are simple inequalities between state potentials, summarized in the following assumption:
Assumption 4.11.
We assume that
  • $\alpha^{d}=0$,
  • all constants $\beta^{d}$, $\alpha^{oe}$, $\beta^{o}$, $\beta^{ne}$ and $\gamma^{ne}$ are non-negative,
  • $\alpha^{oe}\leq\beta^{ne}+\textstyle{\frac{1}{2}}R\leq\beta^{o}+\textstyle{\frac{1}{2}}R$,
  • $\max\{\beta^{d},\beta^{o}\}\leq\gamma^{ne}+\textstyle{\frac{1}{2}}R$,
The bound for $\mathcal{P}_{1}$ is most critical (not surprisingly, as it corresponds to requesting the second item of a pair). To analyze this bound, the needed constraints, besides the state potentials also need to involve the numbers of pairs that are in these states. This gives rise to a non-linear optimization problem that we need to solve.
4.4.1 Analyzing pairs from set $\mathcal{P}_{1}$
The proof of the following lemma is deferred to subsection 4.5.
Lemma 4.12.
There exist parameters $\alpha^{d}$, $\beta^{d}$, $\alpha^{oe}$, $\beta^{o}$, $\beta^{ne}$ and $\gamma^{ne}$, satisfying 4.11, such that for any step $t$, it holds that $\Delta\textsl{{Fpm}}+\sum_{\{x,y\}\in\mathcal{P}_{1}}\Delta\Phi_{xy}\leq R\cdot\sum_{\{x,y\}\in\mathcal{P}_{1}}\Delta w^{xy}$.
4.4.2 Analyzing pairs from set $\mathcal{P}_{2}$
Lemma 4.13.
For any step $t$, it holds that $\sum_{\{x,y\}\in\mathcal{P}_{2}}\Delta\Phi_{xy}\leq R\cdot\sum_{\{x,y\}\in\mathcal{P}_{2}}\Delta w^{xy}$.
Proof 4.14.
Recall that $\mathcal{P}_{2}$ contains all pairs $\{z^{*},y\}$, such that $z^{*}\prec y$. Thus, it is sufficient to show that $\Delta\Phi_{z^{*}y}\leq R\cdot\Delta w^{z^{*}y}$ holds for any such pair $\{z^{*},y\}$. The lemma will then follow by summing over all pairs from $\mathcal{P}_{2}$.
By 4.11, we have
$\alpha^{d}-\alpha^{oe}$ $\leq 0$ (1)
$\max\{\alpha^{d},\alpha^{oe}\}-\beta^{ne}$ $=\alpha^{oe}-\beta^{ne}\leq\textstyle{\frac{1}{2}}R,$ (2)
$\max\{\beta^{d},\beta^{o},\beta^{ne}\}-\gamma^{ne}$ $=\max\{\beta^{d},\beta^{o}\}-\gamma^{ne}\leq\textstyle{\frac{1}{2}}R.$ (3)
We consider two cases depending on the initial mode of the pair $\{z^{*},y\}$. In each case, we upper-bound the potential change on the basis of possible state changes of this pair (cf. Table 2).
  • The initial mode of $\{z^{*},y\}$ is $\alpha$. By Table 2, this mode remains $\alpha$, and thus by subsection 2.2, $\Delta w^{z^{*}y}=0$. Then,
    $\Delta\Phi_{z^{*}y}$ $\leq\max\left\{\alpha^{d}-\alpha^{d},\alpha^{d}-\alpha^{oe}\right\}$ (by Table 2)
    $=0=R\cdot\Delta w^{z^{*}y}.$ (by (1))
  • The initial mode of $\{z^{*},y\}$ is $\beta$ or $\gamma$. By Table 2, its mode changes due to the request to $z^{*}$, and hence, by subsection 2.2, $\Delta w^{z^{*}y}=\textstyle{\frac{1}{2}}$. Then,
    $\Delta\Phi_{z^{*}y}$ $\leq\max\{\alpha^{d}-\alpha^{d},\alpha^{d}-\beta^{d},\alpha^{d}-\alpha^{oe},\alpha^{d}-\beta^{o},$
    $\quad\quad\quad\quad\max\{\alpha^{d},\alpha^{oe}\}-\beta^{ne},$
    $\quad\quad\quad\quad\max\{\beta^{d},\beta^{o},\beta^{ne}\}-\gamma^{ne}\}$ (by Table 2)
    $\leq\max\{-\beta^{d},-\beta^{o},\textstyle{\frac{1}{2}}R,\textstyle{\frac{1}{2}}R\}$ (by $\alpha^{d}=0$, (1), (2) and (3))
    $=\textstyle{\frac{1}{2}}R=R\cdot\Delta w^{z^{*}y}.$
4.4.3 Analyzing pairs from set $\mathcal{P}_{3}$
Lemma 4.15.
For any step $t$, it holds that $\sum_{\{x,y\}\in\mathcal{P}_{3}}\Delta\Phi_{xy}\leq R\cdot\sum_{\{x,y\}\in\mathcal{P}_{3}}\Delta w^{xy}$.
Proof 4.16.
As in the previous lemma, we show that the inequality $\Delta\Phi_{xy}\leq R\cdot\Delta w^{xy}$ holds for any pair $\{x,y\}\in\mathcal{P}_{3}$, i.e., for a pair $\{x,y\}$, such that $x\neq z^{*}$ and $y\neq z^{*}$. The lemma will then follow by summing over all pairs $\{x,y\}\in\mathcal{P}_{3}$.
Possible state transitions of such a pair $\{x,y\}$ are given in Table 3. Hence, such a pair either does not change its state (and then $\Delta\Phi_{xy}=0$) or it changes it from $\beta^{o}$ to $\beta^{ne}$ (and then $\Delta\Phi_{xy}=\beta^{ne}-\beta^{o}\leq 0$ by 4.11). In either case, $\Delta\Phi_{xy}\leq 0\leq R\cdot\Delta w^{xy}$.
4.4.4 Proof of $R$-competitiveness
We now show that the three lemmas above imply that Fpm is $R$-competitive.
Theorem 4.17.
For an appropriate choice of parameters, the competitive ratio of Fpm is at most $R=\textstyle{\frac{1}{8}}(23+\sqrt{17})\leq 3.3904$.
Proof 4.18.
Fix any sequence $\sigma$ consisting of $T$ requests. By summing the guarantees of 4.12, 4.13, and 4.15, we obtain that for any step $t$, it holds that
$\Delta_{t}\textsl{{Fpm}}+\sum_{\{x,y\}\in\mathcal{P}}\Delta_{t}\Phi_{xy}\leq R\cdot\sum_{\{x,y\}\in\mathcal{P}}\Delta_{t}w^{xy}.$
By summing over all steps, observing that the potentials are non-negative and the initial potential is zero (cf. 4.11), we immediately obtain that $\textsl{{Fpm}}(\sigma)\leq R\cdot\sum_{t=1}^{T}\sum_{\{x,y\}\in\mathcal{P}}\Delta w^{xy}\leq R\cdot\textsl{{Opt}}(\sigma).$ The second inequality follows by 2.2.
4.5 Proof of 4.12
Again, we focus on a single step $t$, in which the requested item is denoted $z^{*}$. We let $A^{d}$, $B^{d}$, $A^{oe}$, $B^{o}$, $B^{ne}$, $C^{ne}$ be the number of items $y$ preceding $z^{*}$ such that pairs $\{z^{*},y\}$ have states $\alpha^{d}$, $\beta^{d}$, $\alpha^{oe}$, $\beta^{o}$, $\beta^{ne}$ and $\gamma^{ne}$, respectively. Let
$V\triangleq[A^{d},B^{d},A^{oe},B^{o},B^{ne},C^{ne}].$
Note that $\|V\|_{1}$ is the number of items preceding $z^{*}$, and thus also the access cost Fpm pays for the request. Moreover, $A^{d}+B^{d}$ is the number of items that precede $\theta_{z^{*}}$. We use $\odot$ to denote scalar product (point-wise multiplication) of two vectors.
We define three row vectors $G_{\mathrm{OPT}}$, $G_{\mathrm{PM}}$, and $G_{\mathrm{FM}}$, such that
$\begin{pmatrix}G_{\mathrm{PM}}\\ G_{\mathrm{FM}}\\ G_{\mathrm{OPT}}\end{pmatrix}$ $=\begin{pmatrix}1&1&2&2&2&2\\ 2&2&2&2&2&2\\ \textstyle{\frac{1}{2}}&\textstyle{\frac{1}{2}}&\textstyle{\frac{1}{2}}&\textstyle{\frac{1}{2}}&\textstyle{\frac{1}{2}}&0\end{pmatrix}$
$+\begin{pmatrix}\beta^{ne}&(\gamma^{ne}-\beta^{d})&(\max\{\beta^{d},\beta^{o}\}-\alpha^{oe})&(\alpha^{oe}-\beta^{o})&-\beta^{ne}&-\gamma^{ne}\\ \beta^{d}&-\beta^{d}&(\beta^{d}-\alpha^{oe})&-\beta^{o}&-\beta^{ne}&-\gamma^{ne}\\ 0&0&0&0&0&0\end{pmatrix}$
Expressing costs as vector products.
Recall that to prove 4.12, we need to relate the $\mathcal{P}_{1}$ portion of the amortized cost of Fpm, i.e., $\Delta\textsl{{Fpm}}+\sum_{\{x,y\}\in\mathcal{P}_{1}}\Delta\Phi_{xy}$ and the corresponding portion of the cost of pair-based Opt, i.e., $\sum_{\{x,y\}\in\mathcal{P}_{1}}\Delta w^{xy}$. In the following two lemmas, we show how to express both terms as vector products.
Lemma 4.19.
It holds that $\sum_{y\prec z^{*}}\Delta w^{z^{*}y}=G_{\mathrm{OPT}}\odot V$.
Proof 4.20.
The right-hand side of the lemma relation is equal to ${\frac{1}{2}}(A^{d}+B^{d}+A^{oe}+B^{o}+B^{ne})$. By subsection 4.2, due to request to $z^{*}$:
  • $A^{d}+A^{oe}$ pairs change mode from $\alpha$ to $\beta$, and
  • $B^{d}+B^{o}+B^{ne}$ pairs change mode from $\beta$ to $\gamma$.
By subsection 2.2, each such mode change contributes ${\frac{1}{2}}$ to the left hand side of the lemma equation. Note that $C^{ne}$ pairs of mode $\gamma$ do not change their mode due to the request.
Lemma 4.21.
It holds that $\Delta\textsl{{Fpm}}+\sum_{y\prec z^{*}}\Delta\Phi_{z^{*}y}=G\odot V$, where
  • $G=G_{\mathrm{PM}}$ if Fpm performs a partial move, and
  • $G=G_{\mathrm{FM}}$ if Fpm performs a full move.
Proof 4.22.
First, assume that Fpm performs a partial move. By the definitions of $A^{d}$, $B^{d}$, $A^{oe}$, $B^{o}$, $B^{ne}$, and $C^{ne}$, $\Delta\textsl{{Fpm}}=[1,1,2,2,2,2]\odot V$. By Table 1,
$\sum_{y\prec z^{*}}$ $\Delta\Phi_{z^{*}y}$
$\leq\big{[}\beta^{ne}-\alpha^{d},\gamma^{ne}-\beta^{d},\max\{\beta^{d},\beta^{o},\beta^{ne}\}-\alpha^{oe},\alpha^{oe}-\beta^{o},\alpha^{d}-\beta^{ne},\alpha^{d}-\gamma^{ne}\big{]}\odot V$
$=\left[\beta^{ne},\gamma^{ne}-\beta^{d},\max\{\beta^{d},\beta^{o}\}-\alpha^{oe},\alpha^{oe}-\beta^{o},-\beta^{ne},-\gamma^{ne}\right]\odot V,$
where the second equality follows as $\alpha^{d}=0$ and $\beta^{o}\geq\beta^{ne}$ (by 4.11). Thus, the lemma holds for a partial move.
Next, assume Fpm performs a full move. Then, $\Delta\textsl{{Fpm}}=[2,2,2,2,2,2]\odot V$. Again by Table 1,
$\sum_{y\prec z^{*}}\Delta\Phi_{z^{*}y}$ $=\left[\beta^{d}-\alpha^{d},\alpha^{d}-\beta^{d},\beta^{d}-\alpha^{oe},\alpha^{d}-\beta^{o},\alpha^{d}-\beta^{ne},\alpha^{d}-\gamma^{ne}\right]\odot V$
$=\left[\beta^{d},-\beta^{d},\beta^{d}-\alpha^{oe},-\beta^{o},-\beta^{ne},-\gamma^{ne}\right]\odot V,$
where in the second equality we used $\alpha^{d}=0$ (by 4.11). Thus, the lemma holds for a full move as well.
Recall that Fpm is defined exactly to choose the move that minimizes $\Delta\textsl{{Fpm}}+\sum_{y\prec z^{*}}\Delta\Phi_{z^{*}y}$.
Corollary 4.23.
It holds that $\Delta\textsl{{Fpm}}+\sum_{y\prec z^{*}}\Delta\Phi_{z^{*}y}=\min\{G_{\mathrm{PM}}\odot V,G_{\mathrm{FM}}\odot V\}$.
Finding Parameters.
We may now prove 4.12, restated below for convenience.
See 4.12
Proof 4.24.
We choose the following values of the parameters:
$\alpha^{d}$ $=0$ $\beta^{d}$ $=\textstyle{\frac{1}{16}}(5+3\sqrt{17})\approx 1.086$ $\alpha^{oe}$ $=2$
$\beta^{o}$ $=\textstyle{\frac{1}{16}}(1+7\sqrt{17})\approx 1.866$ $\beta^{ne}$ $=\textstyle{\frac{1}{16}}(9-\sqrt{17})\approx 0.305$ $\gamma^{ne}$ $=2$
It is straightforward to verify that these values satisfy the conditions of 4.11. We note that relation $\alpha^{oe}\leq\beta^{ne}+\textstyle{\frac{1}{2}}R$ holds with equality.
By 4.23, $\Delta\textsl{{Fpm}}+\sum_{\{x,y\}\in\mathcal{P}_{1}}\Delta\Phi_{xy}=\Delta\textsl{{Fpm}}+\sum_{y\prec z^{*}}\Delta\Phi_{z^{*}y}=\min\{G_{\mathrm{PM}}\odot V,G_{\mathrm{FM}}\odot V\}$. On the other hand, by 4.19, $\sum_{\{x,y\}\in\mathcal{P}_{1}}\Delta w^{xy}=\sum_{y\prec z^{*}}\Delta w^{z^{*}y}=G_{\mathrm{OPT}}\odot V$. Hence, it remains to show that $\min\{G_{\mathrm{PM}}\odot V,G_{\mathrm{FM}}\odot V\}\leq R\cdot G_{\mathrm{OPT}}\odot V$.
We observe that
[…]
Let $c=\frac{1}{4}(\sqrt{17}-1)$ and let $G_{\mathrm{COMB}}=c\cdot G_{\mathrm{PM}}+(1-c)\cdot G_{\mathrm{FM}}$. Then,
[…]
Now for any vector $V$,
$\min\{G_{\mathrm{PM}}\odot V,G_{\mathrm{FM}}\odot V\}$ $\leq c\cdot G_{\mathrm{PM}}\odot V+(1-c)\cdot G_{\mathrm{FM}}\odot V$
$=(c\cdot G_{\mathrm{PM}}+(1-c)\cdot G_{\mathrm{FM}})\odot V$
$=G_{\mathrm{COMB}}\odot V=R\cdot G_{\mathrm{OPT}}\odot V,$
completing the proof.
5 Discussion
The question whether there is a $3$-competitive algorithm for List Update with uniform costs remains open. Here we discuss some side results and observations related to this question. Unless otherwise explicitly stated, we assume the partial cost model in this section (that is, accessing location $\ell$ costs $\ell-1$).
While our analysis of Fpm gives an upper bound of $3.3904$, we do not know what is the best ratio achievable with its general approach. For the specific choice of parameters we use in the paper, we verified that Fpm is $3$-competitive for lists of length $3$ (see Appendix D), but not better than $3.04$-competitive for lists of length $5$ (see Appendix E).
For lists of length $3$ there is in fact a much simpler algorithm that uses only a small number of states (at most $6$) to keep track of the past history. Our computational experiments indicate that $3$-competitive algorithms exist for lists with up to $6$ items (see Appendix B).
On the other hand, we also have explored several natural adaptations of algorithms for $\textsf{LUP}_{0}$, showing that none of them has ratio $3$ for $\textsf{LUP}_{1}$. Starting with Algorithm Move-to-Front (Mtf), its $4$-competitiveness follows from its $2$-competitiveness for $\textsf{LUP}_{0}$. The matching lower bound of $4$ is also easy: repeatedly request the last item in the list. Ignoring additive constants, the algorithm pays twice the length of the list at each step, while the static algorithm’s average cost per step is only half of that length. (For partial cost this works even with two items. For full costs, one needs to use arbitrarily long lists.)
The intuition why Mtf performs poorly is that it moves the requested items to front “too quickly”. For the particular strategy in the last paragraph (that always requests the last item), ratio $3$ can be obtained by moving the items to front only every second time they are requested. This algorithm, called Dbit, is a deterministic variant of algorithm Bit from [24] and a special case of algorithm Counter in [24] (with $\ell=2$). Algorithm Dbit was seen as a candidate for optimality in the past (see Footnote 1 in section 1). However, in Appendix C we show, that Dbit is not better than $4$-competitive, and not better than $3.302$-competitive in the full cost model.
The strategies discussed above, including Mtf, Dbit and Counter, either leave the requested item at its current location or move it to front. Generalizing the above bounds, we can show that any strategy limited to such moves cannot achieve ratio better than $3.25$, even for just three items (see subsection C.1). A naive fix would be to, say, move the requested item always half-way towards front. This algorithm is even worse: its competitive ratio is at least $6$ (Appendix C). (Our algorithm Fpm overcomes this difficulty by a much more sophisticated choice of the target location for the requested item.)
Anderson et al. [9] show that the work-function algorithm is $2$-competitive in the $\textsf{LUP}_{0}$ model. Our computer-aided calculation of its performance for short lists indicates that its ratio is $3$ for up to $4$ items, but it is larger than $3$ for $5$ items (see Appendix B).
To shed more light on the $\textsf{LUP}_{1}$ problem, one promising research direction is to examine in depth the concept of projective algorithms, as defined in [8]. The computation of a projective algorithm is uniquely determined by its behavior on two-item instances. As a result, to be $c$-competitive such an algorithm needs to be simultaneously $c$-competitive against each two-item subsequence of the request sequence. This property greatly simplifies competitive analysis, eliminating the need to keep track of the optimum cost for the full instance. Our algorithm Fpm is not projective. We do not know what is the best competitive ratio for $\textsf{LUP}_{1}$ achievable with a projective algorithm.
Appendix A Handling insertions and deletions
In the “dynamic” variant of List Update, insertions and deletions of items are allowed, and cost exactly as much as accessing the item. Crucially, insertions are allowed only at the end of the list; in $\textsf{LUP}_{0}$, the free exchange(s) then allow to immediately move such item to any position for free. As we demonstrate below, any algorithm can be extended to handle insertions while retaining its competitive ratio. [^3] If that ratio depends on the length of the list, i.e., equals $R(\ell)$ for a list of initial (if deletions are allowed) length $\ell$, then $\ell$ equal to the maximum length of the list at any time should be used to determine the ratio. Likewise, any algorithm can be extended to handle deletions but it appears that the effect on its competitive ratio requires further study. Thus, for deletions, we merely demonstrate that Fpm retains its competitive ratio $R$.
Insertions
For any algorithm that does not support insertions (though may support deletions), insertion of $z$ can be treated as an access operation to the item $z$ that was always at the position corresponding to the end of the list immediately after the insertion and was neither accessed nor involved in any exchange before. Specifically, if item $z$ is the $i$-th one to be inserted, then it is treated as having resided exactly $i$ positions after the last item of the initial list. We note in passing that the inserted item $z$ is treated as a new copy of $z$, particularly in the case that $z$ was in the list at any earlier time, either at the very beginning or due to being inserted (possibly multiple times) since then, and that for any prior operation on $z$, the copy which it pertains is trivially determined. Following the insertion, all the actions of the algorithm corresponding to the access operation described are triggered: both internal and those operating on the list, i.e., any exchanges, be they free or paid. are applied as stipulated by the algorithm for the access operation described. Evidently, the cost of handling an insertion this way costs exactly as much as the actually executed access operation.
The same reasoning applies to Opt, pair-based Opt, or the general work function. For pair-based Opt, as $z$ is never accessed prior to insertion, the work function $W^{yz}$ for the pair $\{y,z\}$ remains constant until then, and likewise the pair remains in mode $\alpha$. Generalizing to work functions, it is easy to see that at any time $t$ prior to the insertion of $z$, any state in which $z\prec y$ for some item $y$ in the list at time $t$ is not fundamental.
Now we comment on what this means for Fpm, and why the potential function is unaffected. Initially, $\Theta_{z}=z$ for all items, including those (yet) to be inserted. If $z$ is such an item, then for any item $y\prec z$, we have $\Theta_{y}\preceq y\prec\Theta_{z}=z$, i.e., $\{y,z\}$ pair is in state $\alpha^{d}$, contributing $\alpha^{d}=0$ to the potential. (Note that $y$ may be in the list or itself yet to be inserted.) Any item $y\succ z$ is yet to be inserted if $z$ is yet to be inserted, so we considered all pairs involving some item yet to be inserted. Moreover, since before $z$ is inserted, all the items that are accessed (incl. inserted) or possibly deleted satisfy $y\prec z$ and consequently $\Theta_{y}\prec\Theta_{z}=z$, target cleanups made before the insertion of $z$ neither alter $\Theta_{z}=z$ nor make it a target of another element. Hence, until $z$ is inserted, all the pairs involving $z$ remain in $\alpha^{d}$ state, which contributes 00 to the potential. We note that $\Theta_{z}$ is first changed (to the front of the list) once $z$ is inserted, since target cleanup, movement, and target reset are all performed as the actions corresponding to the access operation.
Deletions
To execute a deletion of $z$, any algorithm simply needs to access $z$ and remove it from the list; in particular, any exchange to be executed immediately after the access that would involve $z$ (which, for all known algorithms, means all the exchanges connected with the access) are obviously not performed. Accounting for the cost is more subtle, because a deleted item $z$ may appear at any position in the list (which may be different for the algorithm and the optimum solution): Even though the cost of deleting $z$ is no different from the cost of accessing it (ignoring aforementioned exchanges that do not take place), once $z$ is deleted, this affects the cost of all subsequent operations on any item(s) $y$ such that $y\succ z$ immediately before $z$’s deletion. Therefore, in terms of costs, this is not equivalent to simply marking $z$ as deleted and ignoring it henceforth. This is why rather than making a general statement, we only argue that Fpm remains $R$-competitive.
To clarify that Fpm is well defined, let us consider work functions first, in order to then focus on their special case, the pair-based Opt. For general work functions, the domain changes (shrinks) after a deletion, since there are fewer elements on the list to be permuted, but the work function definition (and thus the way to compute it) remain unchanged. For the work functions on pairs, involved in cost of pair-based Opt, there are clearly no changes for pairs which do not involve the deleted item. For those which do contain the deleted item $z$, the value of the function is first computed as for a regular access to $z$ and then its value remains frozen forever, since once $z$ is removed from the list, only the other item of the pair remains, never generating any cost. Moreover, we stress that while clearly there is no movement of $z$ by Fpm when it is accessed for deletion and no need to reset $\Theta_{z}$, target cleanup is still (and must be) performed: If any $y\neq z$ had $\Theta_{y}=z$ before the deletion, then after the deletion $\Theta_{y}$ points to $z$’s successor. Clearly, for these to occur, due to the $\Theta_{y}\preceq y$ invariant, $y\succ z$ has to hold, which is why the invariant remains true for $y$ after the deletion.
Hence, when Fpm is to delete $z$, we need to account for the access cost to $z$ (without exchanges) the usual way, taking the change of potential into account. The change of pair potentials is easy to track: For all pairs involving $z$, their state changes to non-existent and stays that way forever since afterwards $z$ is no longer in the list. This means that the pair potential for such pairs changes to 00 regardless of its prior value. Since the prior value of the pair potential is non-negative, the change of potential is at most 00, and as pair-based Opt values never decrease, this implies that 4.13 still holds. As observed before, the target reset may alter some $\Theta_{y}$, changing the state of some pairs that do not involve $z$. But as nothing else changes for those pairs, such changes are still characterized by 4.9 and Table 3, so 4.15 still holds for them.
Thus, to complete the proof that Fpm remains $R$-competitive, it remains to demonstrate that 4.12 continues to hold. Inspecting the proof, one can see that it suffices to show that $G_{\mathrm{del}}\odot V\leq R\cdot G_{\mathrm{OPT}}\odot V$, where $V$ and $G_{\mathrm{OPT}}=\left[\textstyle{\frac{1}{2}},\textstyle{\frac{1}{2}},\textstyle{\frac{1}{2}},\textstyle{\frac{1}{2}},\textstyle{\frac{1}{2}},0\right]$ are as defined before, and $G_{\mathrm{del}}$ is the row vector which captures the cost of deleting $z$ (i.e., accessing $z$, without moving it afterward) and the corresponding change of potential. Defining $G_{\mathrm{del}}$ is easy: accessing $z$ costs $1$ per $\{y,z\}$ pair such that $y\prec z$, and for each such pair the potential changes to 00, so its change is the negative prior value, where those can be read in the first row of Table 1. Thus,
$G_{\mathrm{del}}=\left[1,1,1,1,1,1\right]-\left[0,\beta^{d},\alpha^{oe},\beta^{o},\beta^{ne},\gamma^{ne}\right]$
It is straightforward to observe that for our choice of parameters, $G_{\mathrm{del}}\odot V\leq R\cdot G_{\mathrm{OPT}}\odot V$ does indeed hold: All entries of $G_{\mathrm{del}}$ are at most $1$ because all (subtracted) pairwise potentials are non-negative, moreover $R\geq 3$ and all but the last entry of $G_{\mathrm{OPT}}$ are ${\frac{1}{2}}$. The last entry of $G_{\mathrm{OPT}}$ is 0, but the corresponding one in $G_{\mathrm{del}}$ is $11-\gamma^{ne}=-$ 1.
Appendix B Computational Results
This section is based on results of a computational study of the $\textsf{LUP}_{1}$ problem. Our source code is available at [13].
Note on the accuracy of our results.
Our computational results serve as sources of insight into the $\textsf{LUP}_{1}$ problem, and our code produces graphs which can be checked for correctness, although the size of the graph may be an obstacle. However, it is important to note that we have not verified the correctness of our results by hand nor do we currently include a computer-assisted formal verification of our results.
We therefore list our results as computational results, not as theorems, and advise the reader to be similarly cautious when interpreting them or citing them in future work.
B.1 Upper bound of 3 for list length up to 6
We start by discussing our results of a exhaustive computer search for the optimal competitive ratio for the $\textsf{LUP}_{1}$ model for small list lengths:
Computational result.
There exists a $3$-competitive algorithm for $\textsf{LUP}_{1}$ for list size $n\leq 6$.
Our computational method starts with a fixed competitive ratio $3$ and creates a bipartite graph. In one part (we call it the OPT part), we store all pairs $(\sigma,W)$, where $\sigma$ is the current list of the algorithm and $W$ is a work function for this state. Each vertex in this part has $n$ edges, corresponding to the new request $r$. Each edge leads to a vertex in the ALG part of the graph, which stores a triple ($\sigma$, $W^{\prime}$, $r$), where $W^{\prime}$ is the new work function after including the new request $r$ into the sequence. In turn, each ALG vertex has $n!$ edges, each leading to a different permutation $\sigma^{\prime}$, which we can interpret as the new list of the algorithm after serving the request $r$.
Edges of the graph have an associated cost with them, where the cost of edges leaving $ALG$ correspond to the cost of the algorithm serving the request $r$ and the cost of edges leaving $OPT$ correspond to the increase of the minimum of the work function.
Each vertex in this graph stores a number, which we call the potential of the ALG or OPT vertex. We initialize the potentials to all-zero values and then run an iterative procedure, inspired by [16], which iteratively computes a new minimum potential in each OPT vertex and a new maximum potential in each ALG vertex.
If this procedure terminates, we have a valid potential value for each vertex in the graph, and by restricting our algorithm to pick any tight edge leaving an ALG vertex, we obtain an algorithm that is $3$-competitive. The data generated after this termination are similar in nature to Table 4.
Our computational results correspond to our procedure terminating for list sizes $4\leq n\leq 6$.
For list size $n=6$, it is no longer feasible to create a graph of all reachable work functions, as our computations indicate that even listing all of these would require more memory than our current computational resources allow. We therefore switch to the model of pair-based OPT of Section 2. As explained in that section, a $c$-competitive algorithm for pair-based OPT is also $c$-competitive against regular OPT. However, the space of reachable pair-based work functions is much smaller, so the setting of $n=6$ can be explored computationally as well.
Note that we can also switch the results of $n=4$ and $n=5$ into the pair based setting, which improves the running time. The computational results against the pair-based OPT in those cases are consistent with our results against general OPT.
B.2 Lower bound for the work function algorithm class
We can translate the algorithms for $\textsf{LUP}_{0}$ from [9] to the $\textsf{LUP}_{1}$ model. For input $\sigma$ to $\textsf{LUP}_{1}$ and a list $\pi$, the work function $W^{\sigma}(\pi)$ defines the optimal cost of serving the input $\sigma$ and ending with the list $\pi$. Then, a work function algorithm, given its current state of the list $\mu$ and a new request $r$, it first serves the request and then reorders the list to end in a position which is a minimizer of $\arg\min_{\pi}W^{\sigma,r}(\pi)+d(\pi,\mu)$, where $d(\pi,\mu)$ is the number of pairwise swaps needed to switch from $\mu$ to $\pi$. The minimizer state is not necessarily unique, hence we talk about the work function algorithm class.
Recall that [9] show that some algorithms from the work function algorithm class are $2$-competitive for $\textsf{LUP}_{0}$, which means they are optimal for that setting. In contrast, we show the following:
Computational result.
The competitive ratio of all algorithms in the work function algorithm class in the partial uniform cost model for $\textsf{LUP}_{1}$ is at least $3.1$.
To back up our claim, we present a lower bound instance graph for list length $5$ where all choices of the work function algorithm class are being considered. This graph is equivalent to the graph construction of Section B.1, with the outedges of the ALG vertices restricted only to all valid choices of the work function algorithm class.
However, since the result of an iterative algorithm is that the potential does not stabilize, we provide a heuristical, tailored method how to produce a lower bound graph from the unstable potentials. A result of this code is then a standard lower bound adversary/algorithm graph, with the adversary presenting a single item in each of its vertices and the competitive ratio of any cycle in the graph being at least $3.1$. The source code to generate this graph can be found in our repository [13].
We remark that list length $3$ and $4$ seem to yield a $3$-competitive algorithm for $\textsf{LUP}_{1}$ even from the work function algorithm class.
Appendix C Lower Bounds on MTF and its Modifications
In this section we show the lower bounds greater than $3$ for several natural algorithms, most of which have previously been considered for this problem. All these bounds (unless explicitly stated) hold for the full cost model, and thus also for the partial one.
Deterministic Bit.
Algorithm Deterministic Bit (Dbit): start with all items unmarked. On request to item $r$, if it is marked, move $r$ to front and unmark it. Otherwise, mark $r$. We show that the competitive ratio of Dbit is at least $3.302$.
To this end, consider a list of length $n$ and denote its items as $x_{0},x_{1},\ldots,x_{n-1}$, counting from the front of the list. We divide the list into two parts, call them $A,B$, where part $A$ has length $cn$ and part $B$ has length $(1-c)n$. Value of the parameter $c$ will be determined later. Formally $A=(x_{0},x_{1},x_{2},\ldots,x_{cn-1})$ and $B=(x_{cn},x_{cn+1},\ldots,x_{n-2},x_{n-1})$. To simplify notation, we denote such order of the list as $A,B$. We define request sequence
$\sigma=($ $x_{cn-1},x_{cn-2},\ldots,x_{1},x_{0},$
$x_{n-1},x_{n-1},x_{n-2},x_{n-2},x_{n-3},x_{n-3},\ldots,x_{cn},x_{cn})^{2}.$
In this sequence, all items of $A$ are requested first, starting from the one furthest in the list. Then every item of $B$ is requested twice, also starting from the furthest one. It results in items of $B$ being transported to the front, thus after the first half of $\sigma$ the order of the list of Dbit is $B,A$. Note, that order of items within $B$ does not change.
Then the same sequence of requests is repeated. The difference is that now requesting any item of $A$ results in moving that item to the front. Then items of $B$ are again requested twice, resulting in final list of algorithm being $B,A$. Costs for Dbit are described below.
$\textsl{{Dbit}}(\sigma)=$ $\frac{cn(cn+1)}{2}$ (access to $A$)
$+(1-c)n\cdot(3n-1)$ (access to $B$ and movement)
$+cn\cdot(2n-1)$ (access to $A$ and movement)
$+(1-c)n\cdot(3n-1)$ (access to $B$ and movement)
$=$ $n^{2}\left(6-4c+\frac{c^{2}}{2}\right)+n\left(\frac{3c}{2}-2\right)$ (by simple math)
Now we define a better algorithm for this sequence, which is an upper bound on Opt. It also starts with list $A,B$. After receiving the first batch of requests to the items of $A$, it moves all the items of $B$ before the items of $A$. Then it doesn’t move anymore until the end of request sequence $\sigma$. Costs of such algorithm are described below.
$\textsl{{Opt}}(\sigma)\leq$ $\frac{cn(cn+1)}{2}$ (access to $A$)
$+cn\cdot(1-c)n$ (swap $A$ with $B$)
$+2\cdot\frac{1+(1-c)n}{2}\cdot(1-c)n$ (access to $B$)
$+\frac{(1-c)n+1+n}{2}\cdot cn$ (access to $A$)
$+2\cdot\frac{1+(1-c)n}{2}\cdot(1-c)n$ (access to $B$)
$=$ $n^{2}\left(2-2c+c^{2}\right)+n\left(2-c\right)$ (by simple math)
Note, that the final order of the list in both algorithms is the same and equal to $B,A$. Moreover, every item is requested an even number of times, so no item is marked after $\sigma$. Thus, this request sequence can be repeated infinitely many times, with sets $A,B$ redefined after each sequence.
With long enough list, $n$ becomes insignificant compared to $n^{2}$, therefore the ratio is $\frac{\textsl{{Dbit}}(\sigma)}{\textsl{{Opt}}(\sigma)}\geq\frac{6-4c+\frac{c^{2}}{2}}{2-2c+c^{2}}$. Let $c=\frac{5-\sqrt{13}}{3}$, then $\frac{\textsl{{Dbit}}(\sigma)}{\textsl{{Opt}}(\sigma)}\geq\frac{3+\sqrt{13}}{2}\geq 3.302$.
Deterministic Bit in partial cost model.
Competitive ratio of Dbit in the partial cost model is at least $4$. We will recursively construct a sequence that achieves this ratio for every length of the list $n\geq 2$. For list of length $n$, we denote a sequence of requests as $\sigma_{n}$. For reasons that will become apparent, we want to index elements from the end of the list, namely the algorithm’s list at the start of the sequence is $x_{n-1},x_{n-2},\ldots,x_{1},x_{0}$.
First, assume $n=2$. Set
$\sigma_{2}=x_{1}\,x_{0}\,x_{0}\,x_{1}\,x_{0}\,x_{0}.$
The cost $\textsl{{Dbit}}(\sigma_{2})=8$. Cost $2$ can be achieved for that sequence by starting with the list order $x_{1},x_{0}$ and swapping elements only after the first request. Note that such algorithm would have the same order of elements after serving the whole sequence as algorithm Dbit.
Now let $n>2$. Set
$\sigma_{n}\;=\;x_{n-1}\,\sigma_{n-1}\,x_{n-1}\,\sigma^{\prime}_{n-1}.$
where by $\sigma^{\prime}_{n-1}$ above we mean the sequence obtained from $\sigma_{n-1}$ by permuting the items $x_{n-2},...,x_{0}$ according to their position in the list after executing $\sigma_{n-1}$. Specifically, if $x_{\pi(n-2)}x_{\pi(n-3)}...x_{\pi(0)}$ is the list after executing $\sigma_{n-1}$ then obtain $\sigma^{\prime}_{n-1}$ from $\sigma_{n-1}$ by replacing each item $x_{i}$ by $x_{\pi(i)}$. For example, for $n=3$, we have
$\sigma_{3}\;=\;x_{2},x_{1},x_{0},x_{0},x_{1},x_{0},x_{0},x_{2},x_{0},x_{1},x_{1},x_{0},x_{1},x_{1}.$
The following observation can be proven by straightforward induction. {observation} In $\sigma_{n}$ each item is requested at least twice, and is requested an even number of times. This implies that after executing request sequence $\sigma_{n-1}$ all items $x_{n-2},x_{n-3},...,x_{0}$ will not be marked and the list will have $x_{n-1}$ at the end.
Let $\textsl{{Dbit}}(\sigma_{n})$ be the cost of Dbit on entire input sequence $\sigma_{n}$ and entire list and let $\textsl{{Dbit}}(\sigma_{n-1})$ be the cost of Dbit on $\sigma_{n-1}$ and list of size $n-1$, disregarding element $x_{n-1}$ ($\textsl{{Opt}}(\sigma_{n})$ and $\textsl{{Opt}}(\sigma_{n-1})$ are defined analogously). Thus, we now calculate impact of $x_{n-1}$ on the cost of Dbit.
At the start of the sequence, element $x_{n-1}$ is in front of the list, marked because of the first request. According to Appendix C, during $\sigma_{n-1}$ every element other than $x_{n-1}$ is requested at least twice. For the first pair of such requests $x_{n-1}$ contributes $3$ to the cost (one per each request and one for the swap when the requested element is move to the front). Thus, additional cost contributed by element $x_{n-1}$ for $\sigma_{n-1}$ is at least $3(n-1)$. Then element $x_{n-1}$ is requested and moved to the front with cost $2(n-1)$. After that we apply Appendix C to $\sigma^{\prime}_{n-1}$ so $x_{n-1}$ contributes additional $3(n-1)$. Thus
$\textsl{{Dbit}}(\sigma_{n})\;\geq\;2\cdot\textsl{{Dbit}}(\sigma_{n-1})+8\cdot(n-1).$ (4)
The adversary can serve the first $x_{n-1}$ at cost 00 and then move it immediately to the end, so together with the second request to $x_{n-1}$, it will pay $2(n-1)$ to serve it. So in the final list of the adversary $x_{n-1}$ is at the end, just like in the algorithm’s list. Applying this recursively, we obtain that the adversary’s list after serving is the same as the one for the algorithm. This implies that serving $\sigma_{n-1}$ and $\sigma^{\prime}_{n-1}$ will cost at most $2\textsl{{Opt}}(\sigma_{n-1})$.
$\textsl{{Opt}}(\sigma_{n})\;\leq\;2\cdot\textsl{{Opt}}(\sigma_{n-1})+2\cdot(n-1)$ (5)
Recurrence relations (4) and (5) immediately imply that $\textsl{{Dbit}}(\sigma_{n})/\textsl{{Opt}}(\sigma_{n})\geq 4$.
Half-Move.
Next, we consider the algorithm Half-Move, which moves the requested item to the halfway point between its position and the front of the list. We show that its competitive ratio is at least $6$.
Consider a list of even length $n$. Denote the items as $x_{0},x_{1},\ldots,x_{n-1}$, according to their positions in the initial list of the algorithm. The request sequence consists of $k$ repetitions of requests
$x_{n-1},x_{n-2},x_{n-3},\ldots,x_{\frac{n}{2}}$
for some $k\geq 1$. At the time of the request, the requested item is at the end of the list and it is moved to the middle of the list. Since in the list of even length there are two middle positions, we will assume that the item is always moved to the one further from the front of the list. If it is moved to the other one, the proof only needs to be slightly modified.The cost of any singular request for the algorithm is $n+\frac{n}{2}-1=\frac{3n}{2}-1$. Moreover, after $\frac{n}{2}$ requests, the list returns to its original order. Therefore the total cost of the sequence is $k\cdot\frac{n}{2}\cdot(\frac{3n}{2}-1)$.
It is possible to change the order of the list to $x_{\frac{n}{2}},x_{\frac{n}{2}+1},\ldots,x_{n-1},x_{0},x_{1},x_{2},\ldots,x_{\frac{n}{2}-1}$ before any requests arrive, at cost less than $n^{2}$. Then, with no further movement, total cost of the request sequence is $k\cdot\binom{\frac{n}{2}}{2}$. Thus, the total optimal cost is upper-bounded by $n^{2}+\frac{kn^{2}}{8}$. The ratio approaches $6$ for sufficiently large $n$, as $k\rightarrow\infty$.
C.1 Stay-or-MTF Class
Here we consider a class of online algorithms that never make partial moves. That is, when an item $r$ is requested, the algorithm either does not change the position of $x$ or moves $x$ to the front of the list. We refer to such algorithms as Stay-or-MTF algorithms.
We show that if ${\cal{A}}$ is a Stay-or-MTF algorithm then the competitive ratio of ${\cal{A}}$ is at least $3.25$, even for a list of length $3$.
The adversary strategy is illustrated in the figure below. The list has three items named $a$, $b$ and $c$. The vertices represent states of the game. In each state we assume that the algorithm’s list is $abc$. When an algorithm executes a move, the items are appropriately renamed. Each state is specified by the current offset function. This offset function is equal to the current work function minus the minimum of this work function. At each move, the increase of the minimum of the work function is charged towards the adversary cost, and the offset function is updated. Each offset function that appears in this game is specified by the set of permutations of $a,b,c$ where its value is 00. (Its value at each other list is equal to the minimum swap distance to some offset function with value 00.) To specify this set, we inducate, using braces, which elements are allowed to be swapped. For example, $a{\left\{{bc}\right\}}$ represents two permutations, $abc$ and $acb$, and ${\left\{{abc}\right\}}$ represents all permutations.
The adversary strategy is represented by the transitions between the states. For each state the adversary specifies the request she makes when this state is reached. Notation $r/\xi$ on an arrow means that the request is $x$ and the adversary cost is $\xi$. Then the algorithm decides to make a move. The decision points of the algorithm are indicated by small circles. The type of move is uniquely specified by the cost, specified on the edges from these circles. For example, if the request is on $c$, the algorithm either stays, which costs $2$, or moves to front, which costs $4$.
The numbers next to each state are the potential values. We claim that for each move we have
$\Delta{\cal{A}}+\Delta\Phi\;\geq\;\textstyle{\frac{13}{4}}{\cdot}\Delta\textsl{{Opt}}\,,$
where $\Delta{\cal{A}}$, $\Delta\Phi$ and $\Delta\textsl{{Opt}}$ represent the cost of ${\cal{A}}$, the potential change, and the optimum cost for this move. This inequality can be verified by routine inspection of the transition graph. This shows that the competitive ratio of ${\cal{A}}$ is at least ${\frac{13}{4}}$.
It can be also shown that for $n=3$ ratio ${\frac{13}{4}}$ can be achieved with a Stay-or-MTF algorithm. For $n=4$ we can improve the lower bound to $3.3$, using an approach that is fundamentally similar to the one above, but requiring many more states.
Appendix D A 3-Competitive Algorithm for Lists of Length 3
The Fpm algorithm is $3$-competitive for lists of length 3. This is true even against our lower bound $\sum_{x,y}w^{xy}$ on the optimum cost.
To prove that, we consider possible states of Fpm. There are 15 of them. We assign a potential function $\Psi$ to every state and show that for any state and for any request it holds that $\Delta\textsl{{Fpm}}+\Delta\Psi\leq 3\times\sum_{\{x,y\}\in\mathcal{P}}\Delta w^{xy}$. Note that $\Psi$ is introduced only to simplify the proof and is independent from the algorithm.
The results are shown in Table 4. In each state we assume that the list is abc, and the items are renamed after each transition, if needed. Arrows represent modes: an arrow pointing from an item $x$ towards an item $y$ indicates that the pair $\{x,y\}$ is in mode $\alpha$ (if $x$ is before $y$ in the list) or $\gamma$ (otherwise). Otherwise the pair is in mode $\beta$. Column “MOVE” specifies what Fpm does after the request. Column “DEST.” denotes destination state after movement. State changes corespond to moves of Fpm. Cost of pair-based Opt is in the last column: $P_{OPT}=\sum_{\{x,y\}\in\mathcal{P}}\Delta w^{xy}$.
STATE NUMBER $\Psi$ REQ. MOVE DEST. $\Delta\textsl{{Fpm}}$ $\Delta\Psi$ $P_{OPT}$
[Uncaptioned image] 00 00 $a$ Front 00 00 00 00
$b$ $\Theta_{b}$ $1$ $1$ ${\frac{1}{2}}$ ${\frac{1}{2}}$
$c$ $\Theta_{c}$ $2$ $2$ $1$ $1$
[Uncaptioned image] $1$ ${\frac{1}{2}}$ $a$ Front 00 00 $-\textstyle{\frac{1}{2}}$ ${\frac{1}{2}}$
$b$ Front 00 $2$ $-\textstyle{\frac{1}{2}}$ ${\frac{1}{2}}$
$c$ $\Theta_{c}$ $3$ $2$ $1$ $1$
[Uncaptioned image] $2$ $1$ $a$ Front $4$ 00 $-\textstyle{\frac{1}{2}}$ ${\frac{1}{2}}$
$b$ $\Theta_{b}$ $5$ $1$ $1$ $1$
$c$ Front 00 $4$ $-1$ $1$
[Uncaptioned image] $3$ $\frac{3}{2}$ $a$ Front $4$ 00 $-1$ $1$
$b$ Front $4$ $2$ $-1$ $1$
$c$ Front $4$ $4$ $-1$ $1$
[Uncaptioned image] $4$ ${\frac{1}{2}}$ $a$ Front $4$ 00 00 00
$b$ $\Theta_{b}$ $1$ $1$ 00 $1$
$c$ $\Theta_{c}$ $1$ $3$ 00 $1$
[Uncaptioned image] $5$ $2$ $a$ Front $6$ 00 00 $1$
$b$ Front $4$ $2$ $-\frac{3}{2}$ ${\frac{1}{2}}$
$c$ Front $7$ $4$ $-1$ $1$
[Uncaptioned image] $6$ $2$ $a$ Front $6$ 00 00 00
$b$ $\Theta_{b}$ $1$ $1$ $-\frac{3}{2}$ ${\frac{1}{2}}$
$c$ $\Theta_{c}$ $8$ $3$ 00 $1$
[Uncaptioned image] $7$ $1$ $a$ Front $4$ 00 $-\textstyle{\frac{1}{2}}$ ${\frac{1}{2}}$
$b$ $\Theta_{b}$ $9$ $1$ $1$ $1$
$c$ Front 00 $4$ $-1$ $1$
[Uncaptioned image] $8$ $2$ $a$ Front $10$ 00 $-\textstyle{\frac{1}{2}}$ ${\frac{1}{2}}$
$b$ Front 00 $2$ $-2$ $1$
$c$ $\Theta_{c}$ $11$ $2$ $1$ $1$
[Uncaptioned image] $9$ $2$ $a$ Front 00 00 $-2$ $1$
$b$ Front $10$ $2$ $-\textstyle{\frac{1}{2}}$ ${\frac{1}{2}}$
$c$ $\Theta_{c}$ $12$ $2$ $1$ $1$
[Uncaptioned image] $10$ $\frac{3}{2}$ $a$ Front $10$ 00 00 00
$b$ $\Theta_{b}$ $1$ $1$ $-1$ $1$
$c$ $\Theta_{c}$ $13$ $2$ $1$ $1$
[Uncaptioned image] $11$ $3$ $a$ Front $14$ 00 $-1$ $1$
$b$ Front $7$ $2$ $-2$ $1$
$c$ Front $4$ $4$ $-\frac{5}{2}$ ${\frac{1}{2}}$
[Uncaptioned image] $12$ $3$ $a$ Front $7$ 00 $-2$ $1$
$b$ Front $14$ $2$ $-1$ $1$
$c$ Front $4$ $4$ $-\frac{5}{2}$ ${\frac{1}{2}}$
[Uncaptioned image] $13$ $\frac{5}{2}$ $a$ Front $14$ 00 $-\textstyle{\frac{1}{2}}$ ${\frac{1}{2}}$
$b$ $\Theta_{b}$ $3$ $1$ $-1$ $1$
$c$ Front 00 $4$ $-\frac{5}{2}$ ${\frac{1}{2}}$
[Uncaptioned image] $14$ $2$ $a$ Front $14$ 00 00 00
$b$ $\Theta_{b}$ $8$ $1$ 00 $1$
$c$ $\Theta_{c}$ $1$ $3$ $-\frac{3}{2}$ ${\frac{1}{2}}$
Table 4: Fpm is 3-competitive for lists of length 3
There exists an easier $3$-competitive algorithm for $3$ items. This algorithm uses $6$ states to keep track of the past history. The algorithm is illustrated below as a graph whose vertices represent the current state of the algorithm and edges represent its transitions. Again, items are renamed after each transition so that the list is always abc. The modes of each pair are represented the same way as in Table 4. Each edge is labeled by the requested item, the algorithm’s cost, and the optimum cost (${\frac{1}{2}}$ for each mode change). The algorithm’s move is implied by the cost value. For example, if the request is c and the cost is $3$, it means that the algorithm paid $2$ to access c and then swapped c with the preceding b.
margin: MBa: Here $\Phi$ is not the potential used in the proof. Maybe rename to $\Psi$ to avoid confusion and be consistent with Table 4?
The diagram also shows the potential values $\Phi$ for each state. By routine verification, for each transition we have that the amortized cost (that is, the cost of the algorithm plus the potential change) is at most $3$ times the optimum cost.
Appendix E Lower bound for the Full-Or-Partial-Move algorithm
In this section, we show that for the choice of parematers from 4.12, the competitive ratio of Fpm is at least $3.04$ in the partial cost model.
We construct a sequence that achieves that ratio. That sequence, shown in Table 5, uses only five items, but can be used on any longer list, without requests to any items further than the fifth. Graphs in column “Work function on pairs” represent modes of all the pairs. Arrow from item $x$ to $y$ means that $W^{xy}(xy)+1=W^{xy}(yx)$ (mode $\alpha$ or $\gamma$, depending on the order of the list). Some arrows are omitted for succinctness - the true graph is always the transitive closure of the graph shown.
Note that the final state of the list after the sequence is the same as the state before the sequence, therefore it can be repeated indefinitely. The first state described in the table is accessible from the starting state described in section 3 by requesting item $d$ and then item $a$.
Table 5 compares cost of Fpm to the cost of pair-based Opt defined as $P_{OPT}=\sum_{\{x,y\}\in\mathcal{P}}\Delta w^{xy}$, which for this sequence is equal to $25$. We now show the behaviour of an offline algorithm that achieves this cost. Before the start of the sequence, the algorithm changes its list to $cdeab$. This incurs some additional one time cost, which is negligible if the sequence is repeated sufficiently many times. During the entire sequence, the algorithm performs two moves: after serving the first request, it moves item $c$ to the third position in the list and after ninth request it brings item $c$ back to the front. Thus, the total cost of swaps for the offline algorithm is $4$. It is easy to check that the cost of requests is $21$, so the total cost is $25$.
List of Fpm Work function Req. Move Cost $P_{OPT}$
on pairs of Fpm of Fpm
$a$ $b$ $c$ $d$ $e$ $\theta_{a}$ $\theta_{b}\ \theta_{d}$ $\theta_{c}$ $\theta_{e}$ [Uncaptioned image]
[…]
$a$ $b$ $c$ $d$ $e$ $\theta_{a}\ \theta_{c}$ $\theta_{b}\ \theta_{d}$ $\theta_{e}$ [Uncaptioned image]
[…]
$a\ \ \theta_{a}$ $b$ $c$ $d$ $e$ $\theta_{c}\ \theta_{e}$ $\theta_{b}\ \theta_{d}$ [Uncaptioned image]
[…]
$e$ $a$ $b$ $c$ $d$ $\theta_{e}$ $\theta_{a}\ \theta_{c}$ $\theta_{b}\ \theta_{d}$ [Uncaptioned image]
[…]
$e$ $a$ $d$ $b$ $c$ $\theta_{e}\ \theta_{d}$ $\theta_{a}\ \theta_{c}$ $\theta_{b}$ [Uncaptioned image]
[…]
$c$ $e$ $a$ $d$ $b$ $\theta_{c}$ $\theta_{d}\ \theta_{e}$ $\theta_{a}$ $\theta_{b}$ [Uncaptioned image]
$d$ $\theta_{d}$ 33 11
$c$ $d$ $e$ $a$ $b$ $\theta_{c}\ \theta_{d}$ $\theta_{e}$ $\theta_{a}$ $\theta_{b}$ [Uncaptioned image]
$e$ $\theta_{e}$ 35 12
$c\ \ \theta_{c}$ $d$ $e$ $a$ $b$ $\theta_{d}\ \theta_{e}$ $\theta_{a}$ $\theta_{b}$ [Uncaptioned image]
[…]
$e$ $c$ $d$ $a$ $b$ $\theta_{e}$ $\theta_{c}\ \theta_{d}$ $\theta_{a}$ $\theta_{b}$ [Uncaptioned image]
[…]
$e$ $d$ $c$ $a$ $b$ $\theta_{d}\ \theta_{e}$ $\theta_{c}$ $\theta_{a}$ $\theta_{b}$ [Uncaptioned image]
[…]
$e\ \ \theta_{c}$ $d$ $c$ $a$ $b$ $\theta_{d}\ \theta_{e}$ $\theta_{a}$ $\theta_{b}$ [Uncaptioned image]
[…]
$c$ $e$ $d$ $a$ $b$ $\theta_{c}$ $\theta_{d}\ \theta_{e}$ $\theta_{a}$ $\theta_{b}$ [Uncaptioned image]
[…]
$c$ $e$ $d$ $a$ $b$ $\theta_{b}\ \theta_{c}$ $\theta_{d}\ \theta_{e}$ $\theta_{a}$ [Uncaptioned image]
[…]
$b$ $c$ $e$ $d$ $a$ $\theta_{b}$ $\theta_{c}$ $\theta_{d}\ \theta_{e}$ $\theta_{a}$ [Uncaptioned image]
$d$ $\theta_{d}$ 64 21
$b$ $c$ $d$ $e$ $a$ $\theta_{b}\ \theta_{d}$ $\theta_{c}$ $\theta_{e}$ $\theta_{a}$ [Uncaptioned image]
$a$ $\theta_{a}$ 68 23
$b\ \ \theta_{a}$ $c$ $d$ $e$ $a$ $\theta_{b}\ \theta_{d}$ $\theta_{c}$ $\theta_{e}$ [Uncaptioned image]
$a$ Front 76 25
$a$ $b$ $c$ $d$ $e$ $\theta_{a}$ $\theta_{b}\ \theta_{d}$ $\theta_{c}$ $\theta_{e}$ [Uncaptioned image]
Table 5: Sequence achieving ratio $3.04$ for Fpm. The red arrow indicates the pair in mode $\gamma$.